OpenStudy (anonymous):

why You store less charge on series capacitors than you would on either one of them alone with the same voltage!

OpenStudy (anonymous):

Capacitors i series have the same charge and the voltage will be distributed over them. \[U ={ U_1 + U_2} = {Q \over C_1 }+{Q \over C_2} = Q \left( {1 \over C_1} + {1 \over C_2} \right) = {Q \over C}\]

OpenStudy (anonymous):

for simplicity take 2 capacitors of same capacity c , when they are in series with voltage V then the resultant capacitance will be C =c/ 2 it will store less charge as compared when both capacitors are connected with voltage individually